# Late-Time Evolution of Magnetized Disks in Tidal Disruption Events

### v1

@@ 1 Introduction
Thermal and viscous instabilities are common features in the inner regions of simple accretion disk models around black holes (BHs) (Pringle et al., 1973; Shakura & Sunyaev, 1976; Lightman & Eardley, 1974; Piran, 1978). The inner zones of thin $\alpha$-disk models (Shakura & Sunyaev, 1973) can be describe by three states. Radiation pressure often dominates over the gas pressure in the inner regions, but when the mass accretion rate is very low, radiation pressure becomes inefficient and the disk is stable (Shakura & Sunyaev, 1976; Shen & Matzner, 2014). Alternatively, the disk is also stable at very high accretion rates when advection is the dominant cooling mechanism (Abramowicz et al., 1988; Narayan & Yi, 1995). However, at moderate accretion rates, where radiation pressure and radiative cooling dominate, the disk becomes thermally and viscously unstable. These instabilities are thought to produce global limit-cycle behavior, where the mass accretion rate alternates between the high-accretion-rate, advection-dominated branch and the low-accretion-rate, gas-pressure-dominated state (Honma et al., 1991; Szuszkiewicz & Miller, 1998; Janiuk et al., 2002; Ohsuga, 2006; Shen & Matzner, 2014; Piro & Mockler, 2024).
@@
Despite the theoretical prediction of instabilities in radiatively dominated regions, most observed X-ray binaries (XRBs) and active galactic nuclei (AGNs) appear stable. Evidence of instability has been observed in some cases, such as the ‘heartbeat’ variability in XRBs, a series of bursts with recurrence times ranging from seconds to minutes (Belloni et al., 1997; Taam et al., 1997; Massaro et al., 2010; Altamirano et al., 2011; Bagnoli & in’t Zand, 2015; Maselli et al., 2018), and the long-term variability of changing-look AGNs on timescales of weeks to months (Matt et al., 2013; Bianchi et al., 2005; LaMassa et al., 2015; Ruan et al., 2016; Lawrence, 2018). While these variabilities may be linked to radiation pressure instabilities, most observed XRBs and AGNs remain stable, which challenges the standard $\alpha$-disk predictions (Done et al., 2007).
@@
Another accretion disk phenomenon that exhibits surprising stability is tidal disruption events (TDEs). A TDE occurs when a star approaches sufficiently close to a supermassive black hole (SMBH) and is torn apart by the SMBH’s strong tidal forces (Hills, 1975; Rees, 1988). A portion of the material from the disrupted star is bound to the SMBH, while the other portion is unbound and escapes from the SMBH. Although the early stages of a TDE are complex (Hayasaki et al., 2013; Shiokawa et al., 2015; Hayasaki et al., 2016; Bonnerot & Lu, 2020; Bonnerot & Stone, 2021), in the later stages, the gas dissipates its excess energy and circularizes to form an accretion disk (van Velzen et al., 2019; Steinberg & Stone, 2024).
@@
TDEs have high enough mass accretion rates such that radiation pressure would dominate gas pressure (Ulmer, 1999; Wen et al., 2020). Once the accretion rate falls below the stable advection-dominated state, a 1D $\alpha$-model would predict that the disk should experience global limit-cycle behavior (Shen & Matzner, 2014; Piro & Mockler, 2024). However, observations of late-time TDEs show that the ultraviolet (UV) light curves flatten (Gezari et al., 2015; van Velzen et al., 2019) and generally remain quite constant (Mummery et al., 2024). The observed UV luminosities are too bright to be consistent with the lower, gas-pressure-dominated branch of solutions (van Velzen et al., 2019), which can be orders of magnitude dimmer (Shen & Matzner, 2014; Piro & Mockler, 2024). Therefore, the standard $\alpha$-disk model alone does not explain the late-time TDE observations, indicating that something [^1] While TDE disks are expected to be geometrically complex at early times, with nontrivial tilts (Stone & Loeb, 2012; Franchini et al., 2016), wind losses (Miller, 2015), and global eccentricities (Shiokawa et al., 2015; Zanazzi & Ogilvie, 2020), we are concerned in this paper with very late time observations years post-disruption, at which point the simpler axisymmetric geometry of Shakura-Sunyaev type models should be applicable. prevents thermal and viscous instabilities from developing in these disks.
@@
A common theoretical solution to stabilize the disk involves changing the viscosity parametrization (Sakimoto & Coroniti, 1981; Taam & Lin, 1984; Mummery & Balbus, 2019). However, commonly used viscosity parametrizations are ad hoc and lack physical motivation. It is likely that a more fundamental piece of physics is missing.
@@
Several mechanisms have been proposed to explain the observed stability of most accretion disks, including magnetic fields (Begelman & Pringle, 2007; Oda et al., 2009), viscous fluctuations in a turbulent flow [^2] This mechanism appeals to properties of 3D turbulence. Viscous instabilities have not been extensively studied in 3D MHD simulations because they operate on too long a timescale to easily simulate. However, 1D calculations suggest that thermal and viscous instabilities occur under the same conditions (Piran, 1978), and thermal instabilities have been successfully simulated in 3D MHD (Jiang et al., 2016; Mishra et al., 2022), confirming that they are not merely artifacts of 1D or $\alpha$-viscosity models. (Janiuk & Misra, 2012), delayed pressure responses to stress variations in magnetized disks (Hirose et al., 2009), the “iron opacity bump” (Jiang et al., 2016; Grzędzielski et al., 2017), and strong star-disk collisions (Linial & Metzger, 2024). Among these, stabilization by magnetic fields seems to be the most generic.
@@
Magnetohydrodynamic (MHD) simulations of tidally disrupted stars indicate that the resulting disk’s magnetic field configuration is predominantly toroidal (Guillochon & McCourt, 2017; Bonnerot et al., 2017). These toroidal magnetic fields modify the growth rates of the magnetorotational instability (MRI), which amplifies the toroidal fields. Eventually, the MRI growth rate can be suppressed by magnetic tension, leading to the saturation of the magnetic fields (Pessah & Psaltis, 2005; Begelman & Pringle, 2007). Previous studies debated whether, in the absence of a relatively strong net vertical field, very strong toroidal magnetic flux could remain inside the disk or would escape vertically (Johansen & Levin, 2008; Salvesen et al., 2016). The most recent work on this subject finds that strongly magnetized disks can self-sustain toroidal fields (Squire et al., 2024).
@@
Sufficiently strong magnetic fields can stabilize the disk against thermal and viscous perturbations, provided that magnetic pressure is powerful enough to compete with radiation pressure. The hypothesis that strong magnetic fields stabilize real accretion disks has been substantiated by global, self-consistent 3D radiation MHD simulations (Sądowski, 2016; Jiang et al., 2019; Huang et al., 2023), and more idealized studies (Dexter & Begelman, 2019; Pan et al., 2021; Śniegowska et al., 2023). Magnetic fields comparable in strength to those generated in the aforementioned analytic criteria and MHD simulations were shown by Kaur et al. (2023) to be adequate for stabilizing TDE disks, but so far this has only been demonstrated for steady state disk solutions.
@@
In this paper, we go further and investigate a 1D, time-dependent $\alpha$-disk model where magnetic fields dominate the pressure [^3] We note that while we consider magnetic fields capable of dominating the total pressure budget, these “magnetically elevated” disks generally do not fall into the regime of magnetically arrested (i.e. “MAD;” Narayan et al. 2012) accretion., aiming to study the late-time TDE evolution. Although late-time disks are dim and observationally challenging to study, they are significantly simpler to model compared to the highly complex (Bonnerot & Stone, 2021) early-time TDE behavior. As a result, extracting astrophysical properties, such as the SMBH mass, from early-time behavior remains a considerable challenge. In principle, late-time disks should be simple to model, and therefore useful for estimating astrophysical parameters (Mummery & van Velzen, 2024). However, achieving this goal requires first developing the most accurate model possible.
@@ 2 Magnetized TDE Disks
We develop a 1D, time-dependent thin disk model for a highly magnetized accretion disk, based on the standard Shakura-Sunyaev model (Shakura & Sunyaev, 1973). We use Newtonian gravity, as our main focus is on late-time TDEs when the relevant disk radii, $R$, are large and non-relativistic. The gas in the disk moves around the SMBH with a Keplerian angular frequency $\Omega_{\rm K}=\sqrt{GM_{\bullet}/R^{3}}$, where $M_{\bullet}$ is the SMBH’s mass and $G$ is the gravitational constant. The matter in the disk is steadily accreted into the SMBH; the disk’s surface density $\Sigma(t,R)$ evolves according to the disk diffusion equation:
@@
where $\nu$ is the effective (e.g. turbulent) viscosity. We follow the $\alpha$-disk prescription $\nu=\alpha Hc_{\rm s}$ (Shakura & Sunyaev, 1973). The disk height $H=c_{\rm s}/\Omega_{\rm K}$ is determined from the vertical hydrostatic equilibrium, where $c_{\rm s}=\sqrt{P/\rho}$ is the speed of sound. The density inside the disk is $\rho=\Sigma/H$, and $P$ is the disk pressure.
When the pressure in the disk is radiation-dominated, Eq. 1 describes an unstable disk (Piran, 1978), where the surface density and temperature cycle between high and low accretion states (Shen & Matzner, 2014; Piro & Mockler, 2024). However, as shown in Fig. 1, current late-time TDE observations do not generally display evidence of such instabilities in their late-time light curves, with one possible exception (AT2018dyb; Leloudas et al. 2019). Therefore, in this paper, we focus on magnetized accretion disks, assuming that the disk pressure, $P$, is dominated by magnetic fields, with $P\approx P_{\rm m}=B^{2}/8\pi$, where $B$ is the magnetic field strength.
@@
We note that it is unknown whether angular momentum transport in magnetized disks is primarily driven by local viscous torques, as our use of the $\alpha$-ansatz assumes. Previous research has also examined situations where, instead, large-scale magnetic torques may dominate the angular momentum loss in magnetized disks (Blandford & Payne, 1982; Ferreira & Pelletier, 1993, 1995; Begelman, 2024; Tamilan et al., 2024). Still, these disks may not appear significantly different from standard disks (Begelman, 2024).
To estimate the magnetic field strength, we assume a saturation criterion for the MRI produced when toroidal magnetic fields become “super-thermal” (Pessah & Psaltis, 2005; Begelman & Pringle, 2007). While other saturation criteria have been proposed as well (Oda et al., 2009; Begelman & Armitage, 2023), the one we employ has been roughly validated in recent MHD simulations (Jiang et al., 2019; Mishra et al., 2022; Huang et al., 2023). The MRI is suppressed when the Alfvén speed $v_{\rm A}=\sqrt{P_{\rm m}/\rho}$ exceeds the geometric mean of the Keplerian velocity $v_{\rm K}=\sqrt{GM_{\bullet}/R}$, and the gas sound speed $c_{\rm s,g}=\sqrt{P_{\rm g}/\rho}$, where $P_{\rm g}=\rho\frac{k_{\rm B}T}{\mu m_{\rm p}}$ is the gas pressure. The resulting magnetic pressure is given by:
@@
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant and $\kappa_{\rm{es}}=0.34\,\rm{cm}^{2}\rm{g}^{-1}$ is the Thomson scattering opacity. Electron scattering opacity dominates in the inner regions that interest us and therefore we neglect absorption. Additionally, we neglect advective cooling since, at late times, the disks are thin and sub-Eddington (this neglect would not be justified at early times; conversely, however, advective early-time TDE disks would likely be stable even absent strong magnetic fields; Kaur et al. 2023).
We assume the disk is optically thick, and the emission is isotropic, so the spectral luminosity $L_{\nu}$ at a frequency $\nu$ is given by Planck’s blackbody distribution, $B_{\nu}$, as described in

### v2

@@ 1 Introduction
Thermal and viscous instabilities are common features in the inner regions of simple accretion disk models around black holes (BHs) (Pringle+1973; Shakura&Sunyaev1976; Lightman1974; Piran1978). The inner zones of thin $\alpha$-disk models (Shakura&Sunyaev1973) can be describe by three states. Radiation pressure often dominates over the gas pressure in the inner regions, but when the mass accretion rate is very low, radiation pressure becomes inefficient and the disk is stable (Shakura&Sunyaev1976; Shen&Matzner2014). Alternatively, the disk is also stable at very high accretion rates when advection is the dominant cooling mechanism (Abramowicz1988+; Narayan&Yi1995). However, at moderate accretion rates, where radiation pressure and radiative cooling dominate, the disk becomes thermally and viscously unstable. These instabilities are thought to produce global limit-cycle behavior, where the mass accretion rate alternates between the high-accretion-rate, advection-dominated branch and the low-accretion-rate, gas-pressure-dominated state (Honma+1991; Szuszkiewicz&Miller1998; Janiuk+2002; Ohsuga2006; Shen&Matzner2014; Piro&Mockler2024).
@@
Despite the theoretical prediction of instabilities in radiatively dominated regions, most observed X-ray binaries (XRBs) and active galactic nuclei (AGNs) appear stable. Evidence of instability has been observed in some cases, such as the ‘heartbeat’ variability in XRBs, a series of bursts with recurrence times ranging from seconds to minutes (Belloni+1997; Taam+1997; Massaro+2010; Altamirano+2011; Bagnoli+2015; Maselli+2018), and the long-term variability of changing-look AGNs on timescales of weeks to months (Matt+2013; Bianchi+2005; LaMassa2015; Ruan+2016; Lawrence2018). While these variabilities may be linked to radiation pressure instabilities, most observed XRBs and AGNs remain stable, which challenges the standard $\alpha$-disk predictions (Done+2007).
@@
Another accretion disk phenomenon that exhibits surprising stability is tidal disruption events (TDEs). A TDE occurs when a star approaches sufficiently close to a supermassive black hole (SMBH) and is torn apart by the SMBH’s strong tidal forces (Hills1975; Rees1988). A portion of the material from the disrupted star is bound to the SMBH, while the other portion is unbound and escapes from the SMBH. Although the early stages of a TDE are complex (Hayasaki+2013; Shiokawa+2015; Hayasaki+2016; Bonnerot+2020; Bonnerot&Stone2021), in the later stages, the gas dissipates its excess energy and circularizes to form an accretion disk (vanVelzen+2019; Steinberg&Stone2024).
@@
TDEs have high enough mass accretion rates such that radiation pressure would dominate gas pressure (Ulmer1999; Wen+2020). Once the accretion rate falls below the stable advection-dominated state, a 1D $\alpha$-model would predict that the disk should experience global limit-cycle behavior (Shen&Matzner2014; Piro&Mockler2024). However, observations of late-time TDEs show that the ultraviolet (UV) light curves flatten (Gezari+2015; vanVelzen+2019) and generally remain quite constant (Mummery+2024). The observed UV luminosities are too bright to be consistent with the lower, gas-pressure-dominated branch of solutions (vanVelzen+2019), which can be orders of magnitude dimmer (Shen&Matzner2014; Piro&Mockler2024). Therefore, the standard $\alpha$-disk model alone does not explain the late-time TDE observations, indicating that something [^1] While TDE disks are expected to be geometrically complex at early times, with nontrivial tilts (Stone&Loeb2012; Franchini+2016), wind losses (Miller2015), and global eccentricities (Shiokawa+2015; Zanazzi+20), we are concerned in this paper with very late time observations years post-disruption, at which point the simpler axisymmetric geometry of Shakura-Sunyaev type models should be applicable. prevents thermal and viscous instabilities from developing in these disks.
@@
A common theoretical solution to stabilize the disk involves changing the viscosity parametrization (Sakimoto&Coroniti1981; Taam&Lin1984; Mummery2019). However, commonly used viscosity parametrizations are ad hoc and lack physical motivation. It is likely that a more fundamental piece of physics is missing.
@@
Several mechanisms have been proposed to explain the observed stability of most accretion disks, including magnetic fields (Begelman&Pringle2007; Oda+2009), viscous fluctuations in a turbulent flow [^2] This mechanism appeals to properties of 3D turbulence. Viscous instabilities have not been extensively studied in 3D MHD simulations because they operate on too long a timescale to easily simulate. However, 1D calculations suggest that thermal and viscous instabilities occur under the same conditions (Piran1978), and thermal instabilities have been successfully simulated in 3D MHD (Jiang+2016; Mishra+2022), confirming that they are not merely artifacts of 1D or $\alpha$-viscosity models. (Janiuk&Misra2012), delayed pressure responses to stress variations in magnetized disks (Hirose+2009), the “iron opacity bump” (Jiang+2016; Grzedzielski+2017), and strong star-disk collisions (Linial&Metzger2024). Among these, stabilization by magnetic fields seems to be the most generic.
@@
Magnetohydrodynamic (MHD) simulations of tidally disrupted stars indicate that the resulting disk’s magnetic field configuration is predominantly toroidal (Guillochon&McCourt2017; Bonnerot+2017). These toroidal magnetic fields modify the growth rates of the magnetorotational instability (MRI), which amplifies the toroidal fields. Eventually, the MRI growth rate can be suppressed by magnetic tension, leading to the saturation of the magnetic fields (Pessah&Psaltis2005; Begelman&Pringle2007). Previous studies debated whether, in the absence of a relatively strong net vertical field, very strong toroidal magnetic flux could remain inside the disk or would escape vertically (JohansenLevin2008; Salvesen+2016). The most recent work on this subject finds that strongly magnetized disks can self-sustain toroidal fields (Squire2024).
@@
Sufficiently strong magnetic fields can stabilize the disk against thermal and viscous perturbations, provided that magnetic pressure is powerful enough to compete with radiation pressure. The hypothesis that strong magnetic fields stabilize real accretion disks has been substantiated by global, self-consistent 3D radiation MHD simulations (Sadowski2016; Jiang+2019; Huang+2023), and more idealized studies (Dexter&Begelman2019; Pan+2021; Sniegowska+2023). Magnetic fields comparable in strength to those generated in the aforementioned analytic criteria and MHD simulations were shown by Kaur+2023 to be adequate for stabilizing TDE disks, but so far this has only been demonstrated for steady state disk solutions.
@@
In this paper, we go further and investigate a 1D, time-dependent $\alpha$-disk model where magnetic fields dominate the pressure [^3] We note that while we consider magnetic fields capable of dominating the total pressure budget, these “magnetically elevated” disks generally do not fall into the regime of magnetically arrested (i.e. “MAD;” Narayan+2012) accretion., aiming to study the late-time TDE evolution. Although late-time disks are dim and observationally challenging to study, they are significantly simpler to model compared to the highly complex (Bonnerot&Stone2021) early-time TDE behavior. As a result, extracting astrophysical properties, such as the SMBH mass, from early-time behavior remains a considerable challenge. In principle, late-time disks should be simple to model, and therefore useful for estimating astrophysical parameters (Mummery&vanVelzen2024). However, achieving this goal requires first developing the most accurate model possible.
@@ 2 Magnetized TDE Disks
We develop a 1D, time-dependent thin disk model for a highly magnetized accretion disk, based on the standard Shakura-Sunyaev model (Shakura&Sunyaev1973). We use Newtonian gravity, as our main focus is on late-time TDEs when the relevant disk radii, $R$, are large and non-relativistic. The gas in the disk moves around the SMBH with a Keplerian angular frequency $\Omega_{\rm K}=\sqrt{GM_{\bullet}/R^{3}}$, where $M_{\bullet}$ is the SMBH’s mass and $G$ is the gravitational constant. The matter in the disk is steadily accreted into the SMBH; the disk’s surface density $\Sigma(t,R)$ evolves according to the disk diffusion equation:
@@
where $\nu$ is the effective (e.g. turbulent) viscosity. We follow the $\alpha$-disk prescription $\nu=\alpha Hc_{\rm s}$ (Shakura&Sunyaev1973). The disk height $H=c_{\rm s}/\Omega_{\rm K}$ is determined from the vertical hydrostatic equilibrium, where $c_{\rm s}=\sqrt{P/\rho}$ is the speed of sound. The density inside the disk is $\rho=\Sigma/H$, and $P$ is the disk pressure.
When the pressure in the disk is radiation-dominated, Eq.˜1 describes an unstable disk (Piran1978), where the surface density and temperature cycle between high and low accretion states (Shen&Matzner2014; Piro&Mockler2024). However, as shown in Fig.˜1, current late-time TDE observations do not generally display evidence of such instabilities in their late-time light curves, with one possible exception (AT2018dyb; Leloudas+2019). Therefore, in this paper, we follow Kaur+2023 and focus on magnetized accretion disks, assuming that the disk pressure, $P$, is dominated by magnetic fields, with $P\approx P_{\rm m}=B^{2}/8\pi$, where $B$ is the magnetic field strength.
@@
To estimate the magnetic field strength, we assume a saturation criterion for the MRI produced when toroidal magnetic fields become “super-thermal” (Pessah&Psaltis2005; Begelman&Pringle2007). While other saturation criteria have been proposed as well (Oda+2009; Begelman&Armitage2023), the one we employ has been roughly validated in recent MHD simulations (Jiang+2019; Mishra+2022; Huang+2023). The MRI is suppressed when the Alfvén speed $v_{\rm A}=\sqrt{P_{\rm m}/\rho}$ exceeds the geometric mean of the Keplerian velocity $v_{\rm K}=\sqrt{GM_{\bullet}/R}$, and the gas sound speed $c_{\rm s,g}=\sqrt{P_{\rm g}/\rho}$, where $P_{\rm g}=\rho\frac{k_{\rm B}T}{\mu m_{\rm p}}$ is the gas pressure. The resulting magnetic pressure is given by:
@@
At the high, trans- or super-Eddington accretion rates typical of early-time TDE disks, the Begelman&Pringle2007 ansatz above does not guarantee that $P_{\rm m}\approx P$, as radiation pressure will continue to dominate unless $P_{\rm m}$ is even larger than the value in Eq. 2. However, at the low Eddington ratios relevant for the late-time TDE disks we aim to investigate, Eq. 2 is sufficient to guarantee magnetic stabilization of the disk (and dominance of $P_{\rm m}$ over radiation pressure) at essentially all relevant radii, see e.g. Kaur+2023, Fig. 1.
We note that it is unknown whether angular momentum transport in magnetized disks is primarily driven by local viscous torques, as our use of the $\alpha$-ansatz assumes. Previous research has also examined situations where, instead, large-scale magnetic torques may dominate the angular momentum loss in magnetized disks (Blandford&Payne1982; Ferreira&Pelletier1993; Ferreira&Pelletier1995; Begelman2024; Tamilan+2024). Still, these disks may not appear significantly different from standard disks (Begelman2024).
@@
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant and $\kappa_{\rm{es}}=0.34\,\rm{cm}^{2}\rm{g}^{-1}$ is the Thomson scattering opacity. Electron scattering opacity dominates in the inner regions that interest us and therefore we neglect absorption. Additionally, we neglect advective cooling since, at late times, the disks are thin and sub-Eddington (this neglect would not be justified at early times; conversely, however, advective early-time TDE disks would likely be stable even absent strong magnetic fields; Kaur+2023).
We assume the disk is optically thick, and the emission is isotropic, so the spectral luminosity $L_{f}$ at a frequency $f$ is given by Planck’s blackbody distribution, $B_{f}$, as described in
${\color[rgb]{1,0,0}\textbf{}}L_{f}=4\pi^{2}\int_{R_{\rm in}}^{R_{\rm out}}{\color[rgb]{1,0,0}\textbf{}}B_{f}\left(T_{\rm eff}\right)RdR,$ (4)
where $T_{\rm eff}=T\left(\frac{4}{3\kappa_{\rm es}\Sigma}\right)^{1/4}$ is the effective temperature [^4] For simplicity we ignore the color corrections to emission that cause quasi-thermal accretion disk spectra to deviate from perfect blackbodies (ShimuraTakahara1995). While these corrections can cause $\sim 50\%$ differences between color temperature and effective temperature in X-ray bands, the effect is much smaller at longer wavelengths due to the increased absorption opacity in cooler parts of the disk (DavisElAbd2019)..
The effective viscosity is an approximation to the transport of angular momentum by magnetized turbulence (driven by the MRI). Under the assumption of an $\alpha$-disk prescription, it is given by $\nu=\alpha Hc_{\rm s}=\frac{\alpha c_{s}^{2}}{\Omega_{\rm K}}=\frac{\alpha P}{\Omega_{\rm K}\rho}$, and when the pressure in the disk is dominated by magnetic fields (Eq.˜2), it becomes equal to:
$\nu_{\rm mag}$ $=\nu_{0,\rm{m}}R^{5/7}\Sigma^{2/7}$ (5)
$\nu_{0,\rm{m}}$ $=\left[\frac{27}{32\sigma_{\rm SB}}\alpha^{8}\kappa_{\rm{es}}GM_{\bullet}\left(\frac{k_{\rm B}}{\mu m_{\rm p}}\right)^{4}\right]^{1/7}.$
Because $\nu$ is a function of both $R$ and $\Sigma$ the evolution equation Eq.˜1 becomes a nonlinear diffusion equation. This equation can be separated into a trivial time equation and a nonlinear ordinary differential equation for the radius (see Appendix A, where we show for the first time that a broad class of $\alpha$-viscosity laws can reduce the disk diffusion partial differential equation to the Emden-Fowler ordinary differential equation). However, there is currently no general analytical solution for this equation, necessitating the use of numerical methods.
2.1 Numerical Solution
In this paper, we integrate Eq.˜1 numerically. We set the inner boundary to be the innermost stable circular orbit (ISCO), which for a non-spinning BH equals to $R_{\rm in}=6r_{\rm g}$, where $r_{\rm g}=GM_{\bullet}/c^{2}$ is the gravitational radius of the central BH and $c$ is the speed of light. We assume $\Sigma=0$ at the inner boundary, implies that the viscous torque there is zero, meaning that all the matter that reaches the ISCO is accreted along with its angular momentum. The outer boundary is always chosen to be $R_{\rm out}\gg R_{\rm in}$ such that the surface density $\Sigma$ vanishes at some radius smaller than $R_{\rm out}$ at all times.
In our fiducial models, we use a model in which, initially, half of the mass from the disrupted star has escaped to infinity, while the other half has already circularized to form a ring-like disk at (Strubbe+2009):
$r_{\rm c}=\frac{1+e_{\star}}{\beta}r_{\rm t}=2r_{\rm t}$ (6)
where $e_{\star}$ is the orbital eccentricity of the initially approaching star, and $\beta$ is the ratio of the tidal disruption radius $r_{\rm t}$ to the pericenter distance, and in the final equality we chose $e_{\star}=1$ and $\beta=1$. The tidal radius is (Rees1988):
$r_{\rm t}$ $=\left(\frac{M_{\bullet}}{M_{\star}}\right)^{1/3}R_{\star}$ (7)